OpenStudy (anonymous):

If frac{a}{b} = frac{c}{d}, then prove that frac{a+kc}{b+kd} = frac{a}{b} where k can be all real numbers.

OpenStudy (anonymous):

\[\frac{a}{b} = \frac{c}{d}\]

OpenStudy (anonymous):

show that \[\frac{a+kc}{b+kd} = \frac{a}{b} \]

hero (hero):

I have it

hero (hero):

It wasn't really that hard

hero (hero):

b(a+kc) = a(b+kd) ab + bkc = ab + akd bkc = akd bc = ad c/d =a/b

OpenStudy (anonymous):

i made this myself too, wooo